OpenStudy (anonymous):

This question is on behalf of Aaron7204.

OpenStudy (anonymous):

Combine like terms

OpenStudy (anonymous):

how come aaron is not asking the question himself?

OpenStudy (anonymous):

\[\sqrt{x ^{^{2}}+8x}=4\div \sqrt{3}\]

OpenStudy (anonymous):

Square both sides first

OpenStudy (anonymous):

then solve for zero then factor then solve for x

OpenStudy (anonymous):

\[x ^{2}+8x=16/3\]

OpenStudy (anonymous):

x=-12, ?

OpenStudy (anonymous):

\[3x ^{^{2}}+24x=16\] By cross multiplication

OpenStudy (anonymous):

\[3x ^{2}+24x-16=0\]

OpenStudy (anonymous):

Now solve for x.